• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sulkosky

    Log entry time 22:04:10 on August25,2003

    Entry number 113474

    keyword=Spectrometers Set

    Javier came into the Counting House and fixed the Dipole NMR problem by rebooting the Right Arm IOC crate for the magnets. There was some concern about the reboot ramping the dipole down, but if done properly this will not occur. Below is the procedure to reboot the Right Arm IOC. This does not reboot the Right Q1 because it is on a separate IOC.

    To reset the Right Arm Magnet IOC:

    1.) Go to the middle room of the Counting House and press the green reset button for the Right Arm 'Right Spectrometer Controls VME'.
    Number 16 on the lower panel.

    2.) The read backs on the Hall A Tools GUI will turn white and eventually will come back after a few minutes. Be patient.

    3.) After the read backs come back, Q2 and Q3 will be set to zero. But Q1 and the Dipole will still be at there set values.

    4.) The P0 set parameter will be zero. What ever you do, do not press enter! If you do the dipole will ramp to zero.

    5.) After the read backs are back on Q1 and the Dipole, you should retype the set momentum that you desire and press enter.

    6.) Do not forget to cycle Q2 and Q3, since you are increasing the momentum.

    FIGURE 1